Assume that a country estimates its M1 money supply at $20
million. A broader measure of the money supply, M2, is $50 million.
The country’s gross domestic product is $100 million. Production or
real output for the country is 500,000 units or products.
a. Determine the velocity of money based on the M1 money
supply.
b. Determine the velocity of money based on the M2 money
supply.
c. Determine the average price for the real output.
1
Expert's answer
2011-10-11T12:49:04-0400
M1 =$20 million
M2=$50 million.
GDP= $100 million
Production (real output)= 500,000 units (products)
a. Determine the velocity of money based on the M1 money supply.
b. Determine the velocity of money based on the M2 money supply.
c. Determine the average price for the real output.
The velocity of money (V) is the average frequency with which a unit of money is spent in a specific period of time.
The equation of exchange: MV=PQ
where: M – money supply; P – the average price for the real output; Q –real output. Or PQ is the nominal GDP
a) V1=PQ/M1=$100 million/$20 million=5
b) V2=PQ/M2=$100 million/$50 million=2
c) So we know that VM= $100 million and Q =500,000 units. Where P=VM/Q=100,000,000/500,000=200
Answer: a) 5 b)2 c)200